# Quantum Simulation of Spin-Boson Models with Structured Bath

### v1

@@ Abstract
The spin-boson model, involving a spin interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems. Trapped ions present a natural platform for the quantum simulation of such models; however, previous experiments have been limited to simulating a scenario where a spin is coherently coupled to bosonic modes, which neglects capturing the dissipation of the bath. In our work, we perform quantum simulations of spin-boson models with structured baths using the motional states of trapped ions. We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to the control parameters. Subsequently, we simulate the dynamics of various spin-boson models with spectral densities composed of up to three Lorentzian peaks. The experimental outcomes closely align with theoretical predictions, suggesting the advantage of adding the trapped-ion system’s experimental noise for simulating open quantum systems.
@@
Every quantum system inevitably interacts with the surrounding environment. Such interactions can significantly influence the dynamics of the system. A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al. (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al. (1987), chemical reactions Garg et al. (1985); Gilmore and McKenzie (2005), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018). Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an interesting challenge for quantum simulators Mostame et al. (2012); Leppäkangas et al. (2018); Magazzù et al. (2018); Kim et al. (2022).
@@
Trapped ions provide a highly controllable system of spins and motional modes, and thus are a natural platform for simulating spin-boson models. The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al. (2008) or a few dissipative motional modes Lemmer et al. (2018); Schlawin et al. (2021); MacDonell et al. (2021); Kang et al. (2024). While significant experimental progress has been made in controlling the qubits and motional modes together Jia et al. (2022); Chen et al. (2023), past works in this direction are limited to simulating a spin either coherently coupled to up to two bosonic modes Gorman et al. (2018); MacDonell et al. (2023); Whitlow et al. (2023); Valahu et al. (2023). These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes. Dissipation of the environment determines the system’s long-term behavior, such as reaching a steady state, and therefore is essential to incorporate for accurately modeling open quantum systems. Note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes albeit with limited tunability.
@@
In this work, we perform quantum simulations of spin-boson models with continuous spectral densities using trapped ions and their motional modes. Specifically, we simulate spectral densities composed of several Lorentzian peaks, which is characteristic of baths in light-harvesting protein complexes Olbrich et al. (2011); Chandrasekaran et al. (2015); Lee and Coker (2016); Kim et al. (2018); Tamascelli et al. (2019). Furthermore, we simulate the sub-Ohmic, Ohmic, and super-Ohmic spin-boson models Leggett et al. (1987), which are ubiquitous in various open quantum systems ranging from nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011) and superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023) to molecules in which proton-transfer reactions occur Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). Finally, we simulate the vibrationally-assisted energy transfer (VAET) Gorman et al. (2018) in a simple molecular model where dephasing is added to the vibrational mode. The various spin-state population dynamics, which display behaviors such as damped oscillations, coherent revivals, and relaxation to a steady state, are accurately captured by our quantum-simulation results.
@@
Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters such as laser frequency and phase. Remarkably, adding random frequency offset to the control of trapped ions’ motional modes is directly analogous to the leading source of noise of our experimental setup, namely, motional dephasing Kang et al. (2023a). Thus, motional dephasing noise of the hardware can essentially be used by our quantum simulations, providing accurate results even when the experimental time for simulating the time evolution is comparable to the coherence time of the motional modes. While the quantum simulations presented here are limited to the weak spin-bath coupling regime, our demonstration of adding or using noise Guimarães et al. (2023); Leppäkangas et al. (2023) in analog quantum simulations is an important ingredient for potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al. (2021); Kang et al. (2024).
@@ Results
The spin-boson model is described by a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by ($\hbar$ is set as $1$ for simplicity)
@@
A spin-boson model with structured bath is often characterized by a spectral density $J(\omega)$ composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al. (2018). Specifically, we consider a spin coupled to dephased oscillators (hereafter referred to as the dephased spin-oscillator model). The single-oscillator case is described by the Hamiltonian
@@
where we assume that initially the spin is in the |$\ket{0}$ state and the oscillator is in the thermal state of average phonon number $\bar{n}(\nu)$.
The spectral density of the bath determines the dynamics of the spin up to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup . Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which straightforwardly extends to the model with multiple oscillators. In particular, for the unbiased ($\epsilon=0$) spin-boson model, which is of particular interest Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018), the equilibrium donor-state (|$\ket{0}$) population is $1/2$ for both dephased spin-oscillator models and spin-boson models. Thus, for weak spin-bath coupling, the population dynamics of the two models show an excellent match at all times; the match in short (long) time scale is achieved due to matching spectral densities (equilibrium populations). Unless specified otherwise, we use $\epsilon=0$; results for $\epsilon\neq 0$ can be found in Fig. 5.
A more commonly considered model is a spin coupled to damped oscillators (hereafter referred to as the damped spin-oscillator model); for example, the single-oscillator case is described by the Hamiltonian $\hat{H}_{SO}$ and a pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma(\bar{n}(\nu)+1)}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma\bar{n}(\nu)}\hat{b}^{\dagger}$ Lemmer et al. (2018); Schlawin et al. (2021). For this model, a stronger result holds: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\kappa T$ Tamascelli et al. (2018); sup . However, realizing damped oscillators on trapped ions requires sympathetic cooling on a chain of multiple species of ions Lemmer et al. (2018); Schlawin et al. (2021), which is challenging to perform together with coherent operations that simulate the time evolution. Thus, we instead use dephased oscillators to represent the structured bath, as dephasing can be readily performed using randomized operations on trapped ions as described below.
Experimental results
The dephased spin-oscillator model can be simulated using trapped ions by applying spin-dependent kick (SDK) operations, which induce the coupling between the trapped-ion qubit and the motional mode (see Methods for details). Also, by adding randomness to the control parameters of the SDK operations, we can (i) prepare the thermal state and (ii) simulate the dephasing of the motional mode, by averaging the results over many random trials. The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each with respect to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al. (2017).
@@
First, we demonstrate the tunability of the bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state. Resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details). Figure 1a illustrates the simulated coherent time evolution with respect to $\hat{H}_{SO}$, for various values of $\bar{n}$. The dynamics exhibit coherent oscillations with a slowly fluctuating envelope, where the envelope signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator. This is related to the Jaynes-Cummings collapse and revival with coherent initial states Eberly et al. (1980); our initial thermal state is an ensemble of randomly displaced coherent states. A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to diminished revival amplitude at $t\approx 2\pi/\Delta\times 10.5$. We see a notable agreement between theoretical predictions (solid lines) and experimental results (dots).
@@
We compare our experimental results with the predictions of noise-aware models, which our noisy experiments are expected to be simulating. We use spin-boson models with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma$ [see Eq. (3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments for characterizing imperfect cooling and motional coherence time, respectively. The experimentally measured revival amplitude (Fig. 1c) and fitted $\bar{n}$ (Fig. 1d) match well with the predictions of noise-aware models. This shows that well-characterized experimental noise can be used in determining our control parameters for simulating a target model, and thus is not always a bottleneck for the accuracy of our simulations.
@@
Next, we show the tunability of the dephasing strength, which corresponds to the linewidth of the spin-boson model’s spectral density. Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for detials). This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al. (2020).
Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to critically damped ($\Gamma\approx 0.18\Delta$) to overdamped dynamics as $\Gamma$ is increased. In the underdamped regime, the dephasing strength is sufficiently low such that the amplitude of oscillation revivals is diminished but not entirely vanished. We observe a phase flip occurring at the node of the envelope at $t\approx 2\pi/\Delta\times 5\>(6)$ for $\Gamma=0\>(0.05\Delta)$. Conversely, in the overdamped scenario, strong dephasing diminishes the resonant coupling between the system and the oscillator, leading to oscillations that decay indefinitely without any revivals or phase flips. Thus, after the underdamped oscillations’ revival, the phase of oscillation is opposite between the underdamped and overdamped dynamics, as clearly seen at $t\approx 2\pi/\Delta\times 10$, for example. The experimental data match well with the theoretical predictions of spin-boson models.
@@
The uncertainty of simulated $\bar{n}$ and $\Gamma$ can be reduced by performing a larger number of random trials performed, where each trial uses a fixed set of randomly drawn parameters. In Fig. 1d and e, we perform numerical simulations comprising 2000 random trials and calculate the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, which are obtained by averaging the results over either 20 or 200 trials. In our experiments, due to the control software’s delay in compiling a given set of control parameters, we choose to use 20 random trials, each trial repeated 100 times to obtain the qubit population. The experimental data indeed sit within the standard deviation for 20 trials. In future experiments with faster compilation Dalvi et al. (2023), we expect to perform a larger number of random trials with fewer repetitions each sup .
With the capability of simulating a bath with continuous spectral density, we use multiple oscillators to simulate molecular energy transfer in a model structured bath. We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, where each peak represents the vibrational mode of a molecular system such as light-harvesting protein complexes Olbrich et al. (2011); Chandrasekaran et al. (2015); Lee and Coker (2016); Kim et al. (2018); Tamascelli et al. (2019). We use the electronic coupling strength $\Delta=500$ cm-1 Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
@@
The spin-boson model with $J(\omega)$ in Fig. 2b is simulated using the SDK operations of pulse frequencies shown in Fig. 2a, where the pulse frequency is defined as the motion detuning of laser beams from the carrier transition. The 1st, 3rd, and 5th motional modes (in the order of increasing mode frequency) are used. Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively. The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al. (2010) in the interaction picture Nuomin et al. (2022).
The oscillations in Fig. 2c display repeated suppression and revival in amplitude, where the revival amplitude decreases with time, which is a signature of underdamped oscillations. The second suppression occurs earlier in the 3-mode curve than in the 2-mode curve, as the added Lorentzian peak increases the overall strength of spin-bath coupling. The suppressed oscillations are faster than before or after the suppression, as can be seen from the opposite oscillation phase between the 2-mode and 3-mode curves near $t\approx 650$ fs. Interestingly, the 3-mode experimental data exhibit slower oscillations during the second suppression. Nonetheless, with this exception, the experimental data overall closely align with the theoretical predictions, capturing the fine details of the oscillation envelope.
The experimental time for simulating the time evolution is 1.63 ms (2.56 ms) for the 2-mode (3-mode) experiment, which is in the same order of magnitude as the motional coherence time $5.18(67)$ ms that characterizes motional dephasing, the leading source of errors in our trapped-ion system Kang et al. (2023a). Note that motional dephasing, when modelled as stochastic fluctuations of motional-mode frequencies Wang et al. (2020), simply increases the simulated value of $\Gamma$. A simulation of coherent spin-oscillator model with the same experimental time would have substantial errors due to motional dephasing Kang et al. (2024). However, for simulations of dephased spin-oscillator models, motional dephasing may have negligible effects, as the increase in $\Gamma$ may be small relative to the target value. To improve the accuracy of simulations with longer experimental time, the standard deviation of random frequency offsets may be reduced such that the effects of motional dephasing are compensated. This extends the capability of noisy hardware for simulating open quantum systems by essentially using the noise Guimarães et al. (2023); Leppäkangas et al. (2023). Reliable utilization of noise may require careful characterization, such as spectroscopy of motional dephasing Kang et al. (2023a); Milne et al. (2021); Keller et al. (2021).
We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density is given by
@@
where $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively, and $\omega_{c}$ is the cutoff frequency. This model is widely used to characterize noise in nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011), superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023), and proton-transfer reactions Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or $A$ is increased, which is a subject of longstanding research using analytical tools Leggett et al. (1987); Bulla et al. (2003); Vojta et al. (2005) and classical simulations Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018); Kast and Ankerhold (2013); Nalbach and Thorwart (2013); Otterpohl et al. (2022); Nacke et al. (2023). Adding quantum simulators to the list of methods for studying this model is an interesting task.
We use a sum of Lorentzian spectral densities Meier and Tannor (1999) to approximate the spectral density $J_{\rm Legg}(\omega)$. To find the optimal set of parameters $\kappa$’s, $\Gamma$’s, and $\nu$’s in Eq. (3) that best represent $J_{\rm Legg}(\omega)$, well-established global optimization algorithms Wolz (2022) are used to minimize the cost function, which is defined as the integral of the absolute difference between $J_{\rm Legg}(\omega)$ and the sum of corresponding Lorentzian spectral densities over a given bandwidth sup .
As a first step toward simulating Leggett’s spin-boson models, we consider weak spin-bath coupling ($A=0.1$) and minimize the cost function for a limited bandwidth near resonance $\omega\in[0.9\Delta,1.1\Delta]$ using 2 Lorentzian peaks. Figure 3a shows the obtained sums of Lorentzian peaks (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively. The results for simulating the spin-boson models with these approximated spectral densities are shown in Fig. 3b-d. We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale. The experimental data match well with theoretical predictions, where the deviations in the $s=2$ case result from insufficient number of Trotterization steps sup .
@@
Figure 3e compares the predicted population dynamics of the spin-boson model with spectral density $J_{\rm Legg}(\omega)$ and the corresponding approximated spectral density. While the dynamics match very well for the Ohmic and super-Ohmic cases, the sub-Ohmic case shows a non-negligible deviation. This is because the $J(\omega)$ values are overall larger for the sub-Ohmic bath, so matching the population dynamics requires accurately capturing a larger bandwidth. Indeed, in future experiments with a larger number of motional modes, we expect to match a wider range of spectral density and thus more accurately simulate the sub-Ohmic spin-boson model, as suggested by the 10-mode example in Fig. 4. Note that Ref. Porras et al. (2008) suggests using roughly 100 motional modes to represent the Ohmic spectral density. Using noisy modes allows simulating the spin-boson models much more efficiently in terms of the number of motional modes used.
@@
Finally, we simulate the VAET model Gorman et al. (2018) with nonzero energy detuning $\epsilon$ between the spin states. In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition. Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$.

### v2

@@ Abstract
The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems. Trapped ions present a natural platform for simulating the quantum dynamics of such models, thanks to the presence of both high quality internal qubit states and the motional modes of the ions that can simulate the relevant quantum degrees of freedom. In our work, we extend the previous body of work that focused on coherent coupling of the spins and bosons to perform quantum simulations with structured dissipative baths using the motional states of trapped ions. We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to fully programmable control parameters. Subsequently, we simulate the dynamics of various spin-boson models with noise spectral densities constructed from coupling to several dissipative harmonic oscillator modes. The experimental outcomes closely align with theoretical predictions, indicating successful simulation of open quantum systems using a trapped-ion system.
@@
Interaction of a quantum system with its environment in an open dissipative system can significantly influence the dynamics of the system. A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al. (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al. (1987); Schindler et al. (2022), chemical reactions Garg et al. (1985); Gilmore and McKenzie (2005), collective modes in atomic-photonic systems Strack and Sachdev (2011); Bhaseen et al. (2012), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018). Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an interesting challenge for quantum simulators Mostame et al. (2012); Leppäkangas et al. (2018); Magazzù et al. (2018); Kim et al. (2022).
@@
Trapped ions provide a highly controllable system of spins and motional modes, and thus are a natural platform for simulating spin-boson models. The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al. (2008) or a handful of dissipative motional modes Lemmer et al. (2018); Schlawin et al. (2021); MacDonell et al. (2021); Kang et al. (2024). While significant experimental progress has been made in controlling the qubits and motional modes together Jia et al. (2022); Chen et al. (2023a), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Gorman et al. (2018); MacDonell et al. (2023); Whitlow et al. (2023); Valahu et al. (2023). These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes, that can determine the system’s long-term behavior such as reaching a steady state.
@@
In this work, we perform quantum simulations of spin-boson models with continuous spectral densities using trapped ions and their motional modes. Specifically, we simulate spectral densities composed of several Lorentzian peaks, which is characteristic of baths in light-harvesting protein complexes Olbrich et al. (2011); Chandrasekaran et al. (2015); Lee and Coker (2016); Kim et al. (2018); Tamascelli et al. (2019). Furthermore, we simulate the sub-Ohmic, Ohmic, and super-Ohmic spin-boson models Leggett et al. (1987), which are ubiquitous in various open quantum systems ranging from nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011) and superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023) to molecules in which proton-transfer reactions occur Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). Finally, we simulate the vibrationally-assisted energy transfer (VAET) Gorman et al. (2018) in a simple molecular model where dephasing is added to the vibrational mode. The various spin-state population dynamics, which display behaviors such as damped oscillations, coherent revivals, and relaxation to a steady state, are accurately captured by our quantum-simulation results. We note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
@@
Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters, such as the laser frequency and phase. Remarkably, the impact of these control signals is directly analogous to motional dephasing, the leading source of noise in our experimental setup Kang et al. (2023a). This serves as a lower limit for the level of noise that can be introduced in our quantum simulation process. While the quantum simulations presented here are limited to the weak spin-bath coupling regime, our demonstration of adding or using noise Guimarães et al. (2023); Leppäkangas et al. (2023) in analog quantum simulations is an important ingredient for a potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al. (2021); Kang et al. (2024).
@@ Results
The spin-boson model describes a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by
@@
A spin-boson model with a structured bath is often characterized by a spectral, density $J(\omega)$, composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al. (2018). Specifically, we consider a spin coupled to dephased oscillators (hereafter referred to as the dephased spin-oscillator model). The single-oscillator case is described by the Hamiltonian
@@
where we assume that the oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu)$.
The spectral density of the bath determines the dynamics of the spin to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup . Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which extends straightforwardly to the model with multiple oscillators. Of particular interest is the unbiased ($\epsilon=0$) spin-boson model Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018), where the equilibrium donor-state population is $1/2$ for both dephased spin-oscillator models and spin-boson models. Thus, for weak spin-bath coupling, the population dynamics of the two models show an excellent match at all times; the match in short (long) time scale is achieved due to matching spectral densities (equilibrium populations). Unless specified otherwise, we use $\epsilon=0$; results for $\epsilon\neq 0$ can be found in Fig. 5.
A more commonly considered model is a spin coupled to damped oscillators (hereafter referred to as the damped spin-oscillator model); for example, the single-oscillator case is described by the Hamiltonian $\hat{H}_{SO}$ and a pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma(\bar{n}(\nu)+1)}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma\bar{n}(\nu)}\hat{b}^{\dagger}$ Lemmer et al. (2018); Schlawin et al. (2021). For this model, a stronger result holds: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\kappa T$ Tamascelli et al. (2018); sup . Damped oscillators can be realized in trapped ions using sympathetic cooling on a chain of ions with multiple atomic species or isotopes Lemmer et al. (2018); Schlawin et al. (2021). In this work, we focus on dephased oscillators to represent the structured bath, which is readily performed using programmable randomized control operations on a single species of trapped ions.
Single oscillator model
The dephased spin-oscillator model can be simulated using trapped ions by applying spin-dependent kick (SDK) operations, which induce the coupling between the trapped-ion qubit and the motional mode (see Methods for details). By adding randomness to the control parameters of the SDK operations, we can also prepare the thermal state and simulate the dephasing of the motional mode, by averaging the results over many random trials. The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al. (2017).
@@
We first demonstrate the programmability of the oscillator bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state. Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details). Figure 1a illustrates the simulated coherent time evolution with respect to $\hat{H}_{SO}$, for various values of $\bar{n}$. The dynamics exhibit coherent oscillations with an envelope of slow beat-note that signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator Eberly et al. (1980). Our initial thermal state is prepared as an ensemble of randomly displaced coherent states. A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to reduced beat-note amplitude at $t\approx 2\pi/\Delta\times 10.5$. We see a notable agreement between theoretical predictions (solid lines) and experimental results (dots).
@@
We compare our experimental results with the predictions of noise-aware models, reflecting the noise processes in the experimental setup. We use spin-boson models with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma$ [see Eq. (3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments to characterize imperfect cooling and finite motional coherence time, respectively. The experimentally measured revival amplitude (Fig. 1c) and fitted $\bar{n}$ (Fig. 1d) match well with the predictions of noise-aware models. This shows that well-characterized experimental noise can be used as part of the dissipation model, and thus is not always a bottleneck for the accuracy of our simulations.
@@
Next, we show the tunability of the dephasing strength, which corresponds to the linewidth of the spin-boson model’s spectral density. Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for details). This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al. (2020).
Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$). As the damping is increased beyond the critical point, the beat-note that signifies the excitation transfer from the spin to boson back to spin is suppressed, and there is a monotonic decrease in the oscillation of the donor population in the overdamped regime. We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case (see, for example, the phase difference between the underdamped and overdamped case at $t\approx 2\pi/\Delta\times 10$). These features match very well with the theoretical predictions of spin-boson models.
@@
The uncertainty of $\bar{n}$ and $\Gamma$ realized in the experiment can be reduced by performing a larger number of random trials, each trial using a different set of random parameters. The shaded areas in Fig. 1d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by averaging the results over either 20 (pink) or 200 (blue) random trials. Our experiments are limited to 20 random trials due to the compiling delay required for each set of control parameters, and each trial is repeated 100 times to obtain the expectation value for the donor population. The experimental data is consistent with the standard deviation for 20 trials. We expect to readily increase the number of random trials with fewer repetitions each in future experiments with faster compilation tools Dalvi et al. (2023); sup .
Engineering bath spectral densities
We can use multiple oscillator modes to engineer a target spectral density structure for the oscillator bath. We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, each represented by a motional mode of the ion chain. The spin-boson model with $J(\omega)$ in Fig. 2b is simulated using the SDK operations over three motional modes. Fig. 2a shows the frequency of the laser pulses used to drive the SDK, defined as the detuning of laser beams from the carrier transition frequency. The 1st, 3rd, and 5th motional modes are used, and the detuning of the laser frequency from the motional mode determines the frequency at which the bath spectral density is considered in Figure 2b. Here, we set the electronic coupling strength $\Delta=500$ cm-1 sup ; Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively. The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al. (2010) in the interaction picture Nuomin et al. (2022). The donor population features underdamped coherent oscillations, with contributions from all three modes. The measured population closely follows the theoretical predictions.
The coherence time of the motional modes in our setup of $\sim 5.18(67)$ ms is the leading source of error Wang et al. (2020); Kang et al. (2023a), and is comparable to the experimental time for simulating the time evolution of 2-mode (1.63 ms) and 3-mode (2.56 ms) cases. For these simulations, the contribution of the motional decoherence in the experiment is identical to the random kicks we apply in the spin-oscillator model to simulate the bath, and can be incorporated as the source of dephasing in our simulation with proper calibration Kang et al. (2023a); Milne et al. (2021); Keller et al. (2021).
@@
Simulating the Leggett spin-boson models
We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density of the bath is given by
@@
where $\omega_{c}$ is the cutoff frequency, and $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively. This model is widely used to characterize the noise in nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011), superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023), and proton-transfer reactions Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or $A$ is increased, which has been a subject of longstanding research using analytical tools Leggett et al. (1987); Bulla et al. (2003); Vojta et al. (2005) and classical simulations Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018); Kast and Ankerhold (2013); Nalbach and Thorwart (2013); Otterpohl et al. (2022); Nacke et al. (2023).
We use a sum of Lorentzian lines from motional modes to approximate the spectral density $J_{\rm Legg}(\omega)$ Meier and Tannor (1999). To find the optimal set of parameters $\kappa$’s, $\Gamma$’s, and $\nu$’s in Eq. (3) that best represent $J_{\rm Legg}(\omega)$, well-established global optimization algorithms Wolz (2022) are used to minimize the cost function, defined as the integral of the absolute difference between $J_{\rm Legg}(\omega)$ and the sum of corresponding Lorentzian spectral densities over a given bandwidth sup .
As a first step, we consider weak spin-bath coupling ($A=0.1$) and designed the bath spectrum for a limited bandwidth near resonance $\omega\in[0.9\Delta,1.1\Delta]$ using 2 Lorentzian peaks. Figure 3a shows the obtained spectra from the modes (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively. The results for the donor population dynamics interacting with a bath featuring these approximated spectral densities are shown in Fig. 3b-d. We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale depending on the level of noise in the bath. The experimental data match well with theoretical predictions, where the deviations in the $s=2$ case result from insufficient number of Trotterization steps sup .
@@
Figure 3e compares the predicted population dynamics of the spin-boson model with ideal spectral density $J_{\rm Legg}(\omega)$ and the corresponding approximated spectral density. While the dynamics match very well for the Ohmic and super-Ohmic cases, the sub-Ohmic case shows a non-negligible deviation due to the gap in the two spectral densities within the bandwidth (insets in Fig. 3a). In future experiments with a larger number of motional modes Porras et al. (2008), we expect to match a wider range of spectral densities and thus more accurately simulate the sub-Ohmic spin-boson model, as suggested by the 10-mode numerical example in Fig. 4.
Vibration-assisted energy transfer
@@
Finally, we simulate the VAET model Gorman et al. (2018) with nonzero energy detuning $\epsilon$ between the spin states. Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$. In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition.